# Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity

### 16 Mar 2026 v4

@@ Appendix F Appendix F: Possible connections between flat TDOS and scale–free dynamical phenomena
(V) Consequences for superconducting pairing: In the main text we have shown that the same infrared spectral organization modifies the Cooper channel. The pairing susceptibility takes the form
$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi(\omega),$ (S164)
with
$\Pi(\omega)\sim\int d\lambda\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S165)
For $\rho(\lambda\rightarrow 0)=\rho_{0}$ one obtains
$\Pi(\omega)\sim\frac{1}{|\omega|},$ (S166)
which leads to an algebraically enhanced transition scale
$T_{c}\sim\rho_{0}E_{\rm pair}.$ (S167)
The slow–mode reservoir therefore acts as a dynamical amplifier of intrinsic electronic pairing correlations.
Summary. These observations suggest that the infrared organization of the relaxation spectrum may provide a common dynamical language for several scale–free phenomena observed in correlated systems. Within this perspective the central object is not a single quasiparticle mode but the entire spectrum of collective relaxation processes. Flat infrared TDOS simultaneously produces long–time memory, scale–free lifetime statistics, $1/f$ noise, and algebraic enhancement of superconducting pairing.

### 30 Mar 2026 v5

@@ Appendix E Appendix E: Possible connections between flat TDOS and scale–free dynamical phenomena
(V) Consequences for superconducting pairing: In the main text we have shown that superconducting pairing is governed by the retarded Cooper-channel kernel entering the Thouless criterion. Within the relaxation-spectrum framework the pairing susceptibility is written as
$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi^{R}(\omega),$ (S182)
where the retarded polarization kernel takes the form
$\Pi^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S183)
For a flat infrared TDOS, $\rho(\lambda\rightarrow 0)=\rho_{0}$, the real part of the retarded kernel exhibits logarithmic scaling, $\mathrm{Re}\,\Pi^{R}(\omega)\sim\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right)$. Evaluating at the thermal scale $\omega\sim T$ yields the BCS–like transition condition $1-g_{\rm pair}\,\rho_{0}\ln\frac{\Lambda}{T_{c}}=0$, leading to
$T_{c}\sim\Lambda\exp\!\left(-\frac{1}{g_{\rm pair}\rho_{0}}\right).$ (S184)
It is important to distinguish this retarded response from the fluctuation spectrum of the relaxational modes. The latter is governed by the Keldysh correlator,
$\Pi^{K}(\omega)\sim\int d\lambda\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}\sim\frac{1}{|\omega|},$ (S185)
which reflects the strong infrared enhancement of fluctuations. While this divergence signals the presence of an extensive reservoir of slow modes, it does not directly determine the transition scale.
Thus, in the flat TDOS regime the slow-mode reservoir produces strong infrared fluctuations and long-time memory, but the superconducting transition remains marginal, retaining a BCS-like exponential form. Only when the TDOS becomes infrared-singular, $\rho(\lambda)\sim\lambda^{\alpha}$ with $\alpha<0$, does the retarded kernel itself acquire a power-law divergence, leading to an algebraically enhanced transition scale.
Summary. These results indicate that the infrared organization of the relaxation spectrum provides a unified dynamical framework for a wide range of scale–free phenomena observed in correlated systems.
A flat infrared TDOS generates long–time memory, $K(t)\sim 1/t$, scale–free lifetime distributions, $P(\tau)\sim\tau^{-2}$, and $1/f$ noise, reflecting the presence of a dense continuum of slow relaxation modes.
At the same time, the superconducting instability is governed by the retarded response of this spectrum. In the flat TDOS regime the Cooper channel exhibits logarithmic infrared enhancement, leading to a BCS–like exponential transition scale controlled by the infrared spectral weight. A qualitatively new regime emerges only when the TDOS becomes infrared-singular, where the retarded kernel itself becomes power-law divergent and superconductivity is algebraically enhanced.
Within this perspective, superconductivity and anomalous dynamics arise from a common origin: the infrared spectral structure of the Liouvillian relaxation spectrum, rather than from the exchange of a specific bosonic mediator.
Appendix F Appendix F: Spectral function generated by memory–dominated self–energy
where
Writing $\chi_{R}^{-1}=X(\omega,\mathbf{q})+iY(\omega)$ with
$Y(\omega)$ $=\Gamma\,\mathrm{sgn}(\omega),$ (S191)
one has $\chi_{R}=(X-iY)/(X^{2}+Y^{2})$, hence
which yields
(V) Universal scaling collapse: Define the scaling variable
Using $\Gamma=(\pi/2)\rho_{0}$, Eq. (S192) reduces to the universal form
(VI) Finite–temperature infrared cutoff: At finite temperature the logarithmic growth is regulated by the thermal scale. A minimal implementation consistent with the Matsubara-frequency structure discussed in Appendix D is
$|\omega|\;\rightarrow\;\max\{|\omega|,cT\},\qquad c=O(1),$ (S199)
yielding